# Benchmarking direct and indirect dipolar spin-exchange interactions between two Rydberg atoms

### v1

@@ Abstract
Arrays of Rydberg atoms are now one of the leading platforms for quantum science and technology. They have allowed for several breakthroughs in quantum computing [1, 2, 3, 4], and have proven to be a very versatile tool for analog quantum simulation of spin models [5]. One of the key reasons for this success is the strong, controllable, dipolar interaction between pairs of Rydberg atoms. Measuring accurately the interactions and their coherence time is thus essential for improving the platform and ensuring a faithful implementation of many-body problems of interest [6].
@@
In this article, we experimentally investigate several aspects of the dipolar spin-exchange interaction, which we define as the effective coupling between $\ket{\uparrow,\downarrow}$ and $\ket{\downarrow,\uparrow}$ where $\ket{\uparrow}$ and $\ket{\downarrow}$ are two Rydberg levels. First, we study the influence on the direct spin-exchange of the residual motion of the atoms, that leads to fluctuations in the interatomic distance and thus in the interaction strength: The intrinsic coupling between internal and external degrees of freedom thus acts as a decoherence channel for the internal state. We find a good agreement between our measurements, that use two different methods, and numerical simulations taking into account independently measured experimental imperfections. Second, we extend this study to the indirect spin-exchange that arises when encoding our spin-$1/2$ degrees of freedom in |$\ket{nS}$ and |$\ket{(n+1)S}$ [Fig. 1(c)]. Thanks to the agreement between the simulations and the data, we can analyze the contributions of each imperfection and provide a roadmap for future improvements of the platform. Finally, we demonstrate that, by using an extra light-shifting beam that brings two pair states involving four Zeeman states into resonance, one implements an exotic dipolar coupling term with an unusual angular dependence.
@@ I Experimental setup
Figure 2: Benchmarking a direct dipolar spin-exchange. (a) Two-atom energy levels without (left) and with (right) the dipole-dipole interactions $\hat{V}_{\rm dd}$, showing the initial populations in a spin-exchange experiment. (b) Top: time evolution of the populations of the four pair states, at distance $r=14.9$ µm, $\theta=90$° and with initial trapping frequency $\omega_{\perp}\sim 10$ kHz, for |$\ket{P}=\ket{60P_{3/2},m_{J}=3/2}$: experimental data (points with error bars denoting the standard error on the mean), simulations including (solid line) and excluding (dotted lines) atomic motion. Bottom: illustration of how averaging over many individual oscillations for different interatomic distances leads to damping. (c) Evolution of the quality factor of a spin-exchange experiment with the initial radial trapping frequency $\omega_{\perp}$, which is controlled by the trap depth of the optical tweezers before switching them off. Experimental parameters: $r=17$ µm, $\theta=90$°, |$\ket{P}=\ket{60P_{3/2},m_{J}=1/2}$. The black points (resp. red points) indicate the data with (resp. without) Raman sideband cooling. Lines are simulations taking into account various experimental imperfections: initial positions fluctuations $\sigma_{\perp}$, initial velocity fluctuations $\sigma_{v_{\perp}}$ and decay rates $\Gamma$ from the Rydberg states. The color encodes the radial motional state $\bar{n}_{\perp}$ (and $\bar{n}_{\parallel}\sim 10$). The longitudinal trapping frequency $\omega_{\parallel}$ varies proportionally to $\omega_{\perp}$, but it is not represented since the effect of longitudinal disorder on the Q-factor is negligible. (d,e) Same as (a,b) but for a “two-atom” Ramsey experiment (see text).
@@ II Direct dipolar spin exchange
We first consider the “spin exchange” method [13]. We initialize the atoms in |$\ket{P,S}$ to prepare the symmetric superposition of $\ket{+}$ and $\ket{-}$ [Fig. 2(a)], in which case we expect out-of-phase oscillations at frequency $2J_{SP}$ for the populations of the states |$\ket{P,S}$ and |$\ket{S,P}$. In this method, the oscillation frequency is directly the interaction energy (up to a factor 2) and does not depend on the van der Waals terms; however, it does not give access to the sign of $J_{SP}$. Figure 2(b) shows the experimental results. We observe long-lived oscillations that allow us to extract the interaction energy $|J_{SP}|=2\pi\times(668\pm 1)$ kHz, in agreement with ab-initio calculations using the software PairInteraction [29]. The fit of the oscillations by a sine wave with a Gaussian envelope $\exp(-t^{2}/2\tau^{2})$ yields a damping time $\tau=3.8\pm 0.1$ µs corresponding to a quality factor $Q=\tau|J_{SP}|=16\pm 0.3$.
@@
The numerical simulations work as follows. We first draw initial positions and velocities for the two atoms according to a thermal distribution. We then numerically solve the Lindblad equation, taking into account effects (ii) to (iv). We finally average the results over typically 1000 realizations of the shot-to-shot positional disorder [effect (i)], considering that the atoms are in free flight. The assumption that the atoms are in free flight is only justified at early times, before the van der Waals forces affect the atomic positions (see App. B). The resulting populations are shown as thick solid lines in Fig. 2(b), whereas each realization corresponds to a thin line. Averaging these oscillations, which have slightly different frequencies due to positional disorder, gives rise to the observed damping. Our error model allows us to fully capture the experimental data, including the contrast of the oscillations, using as free parameters $\bar{n}_{\perp}\simeq 0.3$ and $\bar{n}_{\parallel}\simeq 11$ (these values are in reasonable agreement with those obtained by Raman sideband spectroscopy).
Making use of this model, we are now able to dissociate the contributions of the various imperfections to the damping. To do so, we perform another set of measurements using |$\ket{P}=\ket{60P_{3/2},m_{J}=3/2}$ and a distance $r=17$ µm, where we tune the initial position dispersion by changing the depth of the optical tweezers by a factor $\zeta$ before the free flight. The extracted quality factors are plotted in Fig. 2(c) as a function of the initial radial trapping frequency $\omega_{\perp}\propto\sqrt{\zeta}$ (the effect of longitudinal positional disorder on $\Delta r$ being negligible in this regime of parameters). We observe an optimum around $\omega_{\perp}\sim 2\pi\times 50$ kHz, where $Q$ reaches 20. To understand the origin of this maximum, we consider various theoretical situations that include an increasing number of imperfections. First, we consider the ideal case of trapped Rydberg atoms ($\sigma_{v_{i}}=0$) in their motional ground state ($\bar{n}_{\perp}=\bar{n}_{\parallel}=0$), with infinite lifetimes. Here, the solution can be worked out analytically (see App. A): the damping envelope is a Gaussian with a quality factor $Q=r/(6\Delta r)$. In this case, the damping fundamentally comes from the dephasing of the spatial components of the atomic wavefunction, and it can be improved by reducing the positional disorder as $\sigma_{\perp}\propto\omega_{\perp}^{-1/2}\propto\zeta^{-1/4}$ [dotted line in Fig. 2(c)]. Second, we add the finite Rydberg lifetimes and numerically obtain a reduced quality factor (dashed line), that still increases with $\omega_{\perp}$ (and eventually saturates when the lifetime becomes the dominant contribution). Third, we consider atoms in free flight with a distribution of velocities $\sigma_{v\perp}=\omega_{\perp}\sigma_{\perp}$. The finite velocities result in an increased positional disorder, which overcomes the initial positional disorder for a given value of $\omega_{\perp}$, and thus qualitatively explains the observed optimum [blue solid line in Fig. 2(c)]. Finally, to obtain a quantitative agreement with the data, we need to include the finite temperature and obtain the best agreement for $\bar{n}_{\perp}=1.1\pm 0.3$ [gray solid line in Fig. 2(c)], slightly higher than previously quoted for this new set of measurement.
@@
Another method for measuring the dipole-dipole interactions is an adaptation of a Ramsey sequence to the case of two atoms. Starting from |$\ket{S,S}$ after the global Rydberg excitation, we apply a global microwave $\pi/2$ pulse to prepare the state $\ket{\rightarrow,\rightarrow}$ where |(|$\ket{\rightarrow}\equiv\left(\ket{S}+\ket{P}\right)/\sqrt{2}$ corresponds to a spin along $x$. We let the system evolve for a time $t$ and apply another global microwave $\pi/2$ pulse to measure in the $x$ basis. Unlike above, this sequence does not require local addressing, which can induce additional decoherence (reduced lifetime for the addressed state |$\ket{S}$, momentum kicks due to the ponderomotive force). However, contrary to a spin-exchange experiment, the initial state gets decomposed over three eigenstates of $H$, each one with a different eigenenergy: |$\ket{\rightarrow,\rightarrow}=\frac{1}{\sqrt{2}}\left[\ket{+}+\frac{1}{\sqrt{2}}\left(\ket{S,S}+\ket{P,P}\right)\right]$ [Fig. 2(d)]. In the absence of a careful phase compensation, this would lead to a beating instead of a simple oscillation. This beating can be compensated by a simple microwave detuning $\delta=(V_{PP}-V_{SS})/2$, such that the accumulated phase for the two states |$\ket{S,S}$ and |$\ket{P,P}$ is the same  [^3] An alternative method would be to use a spin echo technique, such as done in [47].. After the second $\pi/2$-pulse, one expects an oscillation between the states |$\ket{S,S}$ and |$\ket{P,P}$ with an angular frequency as $\tilde{J}_{SP}\equiv J_{SP}+V_{SP}-(V_{SS}+V_{PP})/2$. As a result, this method informs about the relative sign between $J_{SP}$ and van der Waals terms.
@@
Experimentally, we optimize the microwave detuning $\delta$ by maximizing the contrast of the oscillation, resulting in the curve shown in Fig. 2(e). We obtain $|\tilde{J}_{SP}|=2\pi\times(635\pm 1)$ kHz, in agreement with the calculated interaction energies $J_{SP}=-2\pi\times 675$ kHz, $V_{SS}=2\pi\times 12$ kHz, $V_{SP}=2\pi\times 17$ kHz and $V_{PP}=-2\pi\times 52$ kHz. The initial contrast and the damping of the oscillation are again well explained by simulations with the same experimental imperfections as above. Here, the damping arises mainly from positional disorder; Rydberg lifetimes contribute to a lesser extent  [^4] In addition, a slight overcompensation of the detuning, which was experimentally set to $\delta=-2\pi\times 40$ kHz whereas the ideal compensation should happen for $\delta=-2\pi\times 32$ kHz according to the simulations, leads to a small rise up of the probabilities of the states |$\ket{S,P}$ and |$\ket{P,S}$ at long times due to the above-mentioned beating phenomenon, mimicking additional damping.. Compared with the case of the spin exchange with the same experimental parameters [Fig. 2(b)], the damping time is twice as long, but the oscillation frequency is also twice as small, so that the quality factor is approximately the same ($Q=13.8\pm 0.2$).
@@ IV A spin exchange made resonant by local addressing
where $\theta$ and $\phi$ are the spherical coordinates of the interatomic axis $\mathbf{e}_{r}=\sin(\theta)\cos(\phi)\mathbf{e}_{z}+\sin(\theta)\sin(\phi)\mathbf{e}_{x}+\cos(\theta)\mathbf{e}_{y}$. In Eq. (3), the first line corresponds to the well-known $1-3\cos^{2}(\theta)$ angular dependence measured in Fig. 2(e), whereas the $J_{SP,P^{\prime}S^{\prime}}$ coupling originates from the second line ($d_{1}^{+}d_{2}^{0}$ term), with an expected $\sin(2\theta)$ angular dependence, together with a phase factor depending on $\phi$. Contributions of this term have already been observed in Rydberg dressing experiments [40].
@@ V Outlook
In the experiments reported here, the motion of the atoms leads to dephasing of the spin-exchange oscillations but, since we restrict ourselves to relatively short times, the dipolar interaction does not influence the motion of the atoms (as discussed in App. B). However, going one step further, one can also take into account the mechanical effect of van der Waals or dipolar interactions on the motion of the atoms. This gives rise to a spin-motion coupling, observed for instance in [44], that yields entanglement between internal and external degrees of freedom. From the perspective of quantum simulation of spin models, this is yet another decoherence channel; however, this spin-motion coupling may also open interesting perspectives. For example, considering that the relative atomic motion is described by a one dimensional harmonic oscillator with annihilation and creation operators $\hat{a},\hat{a}^{\dagger}$: $\hat{r}=r_{0}+\Delta x(\hat{a}+\hat{a}^{\dagger})$, the Taylor expansion of Eq. (1) contains a term of the form $(|J|\Delta x/r)(\hat{a}+\hat{a}^{\dagger})(\hat{\sigma}_{1}^{+}\hat{\sigma}_{2}^{-}+\hat{\sigma}_{1}^{-}\hat{\sigma}_{2}^{+})$ (with $\hat{\sigma}^{\pm}$ the pseudo-spin operators), which could be tuned. Spin-motion coupling has also be suggested to implement GKP codes [45] for quantum error correction.
@@ Acknowledgements.
Appendix A Effect of positional disorder onto the damping of a spin exchange oscillation

### v2

@@ Abstract
Arrays of Rydberg atoms are now one of the leading platforms for quantum science and technology. They have allowed for several breakthroughs in quantum computing [1, 2, 3, 4], and have proven to be a very versatile tool for analog quantum simulation of spin models [5]. One of the key reasons for this success is the strong, controllable, dipolar interaction between pairs of Rydberg atoms. Measuring accurately the interactions and their coherence time is thus essential for improving the platform and ensuring a faithful implementation of many-body problems of interest [6]. Such benchmarks have been realized in the context of ultrafast Rydberg interactions [7], Rydberg-dressing [8] and polar molecules [9, 10, 11, 12] and pointed to the crucial role of positional disorder as a source of dephasing in quench experiments. Here, we aim at performing this benchmark for typical experimental parameters used in current intermediate-scale quantum simulations with Rydberg arrays [13, 14, 15, 16].
@@
In this article, we experimentally investigate several aspects of the dipolar spin-exchange interaction, which we define as the effective coupling between $\ket{\uparrow,\downarrow}$ and $\ket{\downarrow,\uparrow}$ where $\ket{\uparrow}$ and $\ket{\downarrow}$ are two Rydberg levels. First, we study the influence on the direct spin-exchange of the residual motion of the atoms, that leads to fluctuations in the interatomic distance and thus in the interaction strength. We find a good agreement between our measurements, that use two different methods, and numerical simulations taking into account independently measured experimental imperfections. Second, we extend this study to the indirect spin-exchange that arises when encoding our spin-$1/2$ degrees of freedom in |$\ket{nS}$ and |$\ket{(n+1)S}$ [Fig. 1(c)]. Thanks to the agreement between the simulations and the data, we can analyze the contributions of each imperfection and provide a roadmap for future improvements of the platform. Finally, we demonstrate that, by using an extra light-shifting beam that brings two pair states involving four Zeeman states into resonance, one implements an exotic dipolar coupling term with an unusual angular dependence.
@@ I Experimental setup
Figure 2: Benchmarking a direct dipolar spin-exchange. (a) Two-atom energy levels without (left) and with (right) the dipole-dipole interactions $\hat{V}_{\rm dd}$, showing the initial populations in a spin-exchange experiment (pink disks) and in a two-atom Ramsey experiment (blue disks). (b) Top: time evolution of the populations of the four pair states, at distance $r=14.9$ µm, $\theta=90$° and with initial trapping frequency $\omega_{\perp}\sim 2\pi\times 10$ kHz, for |$\ket{P}=\ket{60P_{3/2},m_{J}=3/2}$: experimental data (points with error bars denoting the standard error on the mean), simulations including (solid line) and excluding (dotted lines) atomic motion. Bottom: illustration of how averaging over many individual oscillations for different interatomic distances leads to damping. (c) Evolution of the quality factor of a spin-exchange experiment with the initial radial trapping frequency $\omega_{\perp}$, which is controlled by the trap depth of the optical tweezers before switching them off (the atoms being in free flight during the Rydberg sequence). Experimental parameters: $r=17$ µm, $\theta=90$°, |$\ket{P}=\ket{60P_{3/2},m_{J}=1/2}$. The black points (resp. red points) indicate the data with (resp. without) Raman sideband cooling. Lines are classical simulations taking into account various experimental imperfections (see text). The color encodes the radial motional state $\bar{n}_{\perp}$ (and $\bar{n}_{\parallel}\sim 10$). The longitudinal trapping frequency $\omega_{\parallel}$ varies proportionally to $\omega_{\perp}$, but it is not represented since the effect of longitudinal disorder on the quality factor is negligible. (d) Same as (b) but for a “two-atom” Ramsey experiment (see text).
@@ II Direct dipolar spin exchange
We first consider the “spin exchange” method [23]. We initialize the atoms in |$\ket{P,S}$ to prepare the symmetric superposition of $\ket{+}$ and $\ket{-}$ [pink disks in Fig. 2(a)], in which case we expect out-of-phase oscillations at frequency $2J_{SP}$ for the populations of the states |$\ket{P,S}$ and |$\ket{S,P}$. In this method, the oscillation frequency is directly the interaction energy (up to a factor 2) and does not depend on the van der Waals terms; however, it does not give access to the sign of $J_{SP}$. Figure 2(b) shows the experimental results. We observe long-lived oscillations that allow us to extract the interaction energy $|J_{SP}|=2\pi\times(668\pm 1)$ kHz, in agreement with ab-initio calculations using the software PairInteraction [35]. The fit of the oscillations by a sine wave with a Gaussian envelope $\exp(-t^{2}/2\tau^{2})$ yields a damping time $\tau=3.8\pm 0.1$ µs corresponding to a quality factor $Q=\tau|J_{SP}|=16\pm 0.4$.
@@
The numerical simulations work as follows. We first draw initial positions and velocities for the two atoms according to a thermal distribution. We then numerically solve the Lindblad equation, taking into account effects (ii) to (iv). We finally average the results over typically 1000 realizations of the shot-to-shot positional disorder [effect (i)], considering that the atoms are in free flight. The assumption that the atoms are in free flight is only justified at early times, before the van der Waals forces affect the atomic positions (see App. A). The resulting populations are shown as thick solid lines in Fig. 2(b), whereas each realization corresponds to a thin line. Averaging these oscillations, which have slightly different frequencies due to positional disorder, gives rise to the observed damping. Our error model allows us to fully capture the experimental data, including the contrast of the oscillations, using as free parameters $\bar{n}_{\perp}\simeq 0.3$ and $\bar{n}_{\parallel}\simeq 11$ (these values are in reasonable agreement with those obtained by Raman sideband spectroscopy).
To further test the influence of atomic motion onto the damping of the spin-exchange, we perform another set of measurements using |$\ket{P}=\ket{60P_{3/2},m_{J}=3/2}$ and a distance $r=17$ µm, where we scan the initial position and velocity dispersions by changing the depth of the optical tweezers by a factor $\zeta$ before the free flight. The extracted quality factors are plotted in Fig. 2(c) as a function of the initial radial trapping frequency $\omega_{\perp}\propto\sqrt{\zeta}$ (the effect of longitudinal positional disorder on $\Delta r$ being negligible in this regime of parameters). We observe an optimum around $\omega_{\perp}\sim 2\pi\times 50$ kHz, where $Q$ reaches 20. Numerical simulations with the same model [solid lines in Fig. 2(c)] predict a similar behavior; we obtain the best agreement for $\bar{n}_{\perp}=1.1\pm 0.3$ [gray solid line], slightly higher than previously quoted for this new set of measurements. The optimum originates from a trade-off between the initial position dispersion $\sigma_{\perp}\propto 1/\sqrt{\omega_{\perp}}$ and the initial velocity dispersion $\sigma_{v_{\perp}}\propto\sqrt{\omega_{\perp}}$: at low trapping frequencies, $\sigma_{\perp}$ is the major contribution to the damping, whereas at high trapping frequencies $\sigma_{v_{\perp}}$ dominates. In App. B, we design a model to better understand this behavior; we also study the case of trapped Rydberg atoms and show that it could significantly improve the damping for trapping frequencies $\omega_{\perp}>2\pi\times 30$ kHz.
@@
Another method for measuring the dipole-dipole interactions is an adaptation of a Ramsey sequence to the case of two atoms. Starting from |$\ket{S,S}$ after the global Rydberg excitation, we apply a global microwave $\pi/2$ pulse to prepare the state $\ket{\rightarrow,\rightarrow}$ where |(|$\ket{\rightarrow}\equiv\left(\ket{S}+\ket{P}\right)/\sqrt{2}$ corresponds to a spin along $x$. We let the system evolve for a time $t$ and apply another global microwave $\pi/2$ pulse to measure in the $x$ basis. Unlike above, this sequence does not require local addressing, which can induce additional decoherence (reduced lifetime for the addressed state |$\ket{S}$, momentum kicks due to the ponderomotive force). However, contrary to a spin-exchange experiment, the initial state gets decomposed over three eigenstates of $H$, each one with a different eigenenergy: |$\ket{\rightarrow,\rightarrow}=\frac{1}{\sqrt{2}}\left[\ket{+}+\frac{1}{\sqrt{2}}\left(\ket{S,S}+\ket{P,P}\right)\right]$ [blue disks in Fig. 2(a)]. In the absence of a careful phase compensation, this would lead to a beating instead of a simple oscillation. This beating can be compensated by a simple microwave detuning $\delta=(V_{PP}-V_{SS})/2$, such that the accumulated phase for the two states |$\ket{S,S}$ and |$\ket{P,P}$ is the same. An alternative method would be to use a dynamical decoupling technique, such as done in [10, 9]. After the second $\pi/2$-pulse, one expects an oscillation between the states |$\ket{S,S}$ and |$\ket{P,P}$ with an angular frequency as $\tilde{J}_{SP}\equiv J_{SP}+V_{SP}-(V_{SS}+V_{PP})/2$. As a result, this method informs about the relative sign between $J_{SP}$ and van der Waals terms.
@@
Experimentally, we optimize the microwave detuning $\delta$ by maximizing the contrast of the oscillation, resulting in the curve shown in Fig. 2(d). We obtain $|\tilde{J}_{SP}|=2\pi\times(635\pm 1)$ kHz, in agreement with the calculated interaction energies $J_{SP}=-2\pi\times 675$ kHz, $V_{SS}=2\pi\times 12$ kHz, $V_{SP}=2\pi\times 17$ kHz and $V_{PP}=-2\pi\times 52$ kHz. The initial contrast and the damping of the oscillation are again well explained by simulations with the same experimental imperfections as above. Here, the damping arises mainly from positional disorder; Rydberg lifetimes contribute to a lesser extent  [^3] In addition, a slight overcompensation of the detuning, which was experimentally set to $\delta=-2\pi\times 40$ kHz whereas the ideal compensation should happen for $\delta=-2\pi\times 32$ kHz according to the simulations, leads to a small rise up of the probabilities of the states |$\ket{S,P}$ and |$\ket{P,S}$ at long times due to the above-mentioned beating phenomenon, mimicking additional damping.. Compared with the case of the spin exchange with the same experimental parameters [Fig. 2(b)], the damping time is twice as long, but the oscillation frequency is also twice as small, so that the quality factor is approximately the same ($Q=13.8\pm 0.2$).
@@ IV A spin exchange made resonant by local addressing
where $\theta$ and $\phi$ are the spherical coordinates of the interatomic axis $\mathbf{e}_{r}=\sin(\theta)\cos(\phi)\mathbf{e}_{z}+\sin(\theta)\sin(\phi)\mathbf{e}_{x}+\cos(\theta)\mathbf{e}_{y}$. In Eq. (3), the first line corresponds to the well-known $1-3\cos^{2}(\theta)$ angular dependence measured in Fig. 3(b), whereas the $J_{SP,P^{\prime}S^{\prime}}$ coupling originates from the second line ($d_{1}^{+}d_{2}^{0}$ term), with an expected $\sin(2\theta)$ angular dependence, together with a phase factor depending on $\phi$. Contributions of this term have already been observed in Rydberg dressing experiments [45].
@@ V Outlook
In the experiments reported here, the motion of the atoms leads to dephasing of the spin-exchange oscillations but, since we restrict ourselves to relatively short times, the dipolar interaction does not influence the motion of the atoms (as discussed in App. A). However, going one step further, one can also take into account the mechanical effect of van der Waals or dipolar interactions on the motion of the atoms. This gives rise to a spin-motion coupling, observed for instance in [7, 49], that yields entanglement between internal and external degrees of freedom. From the perspective of quantum simulation of spin models, this is yet another decoherence channel; however, this spin-motion coupling may also open interesting perspectives. For example, considering that the relative atomic motion is described by a one dimensional harmonic oscillator with annihilation and creation operators $\hat{a},\hat{a}^{\dagger}$: $\hat{r}=r_{0}+\Delta x(\hat{a}+\hat{a}^{\dagger})$, the Taylor expansion of Eq. (1) contains a term of the form $(|J|\Delta x/r)(\hat{a}+\hat{a}^{\dagger})(\hat{\sigma}_{1}^{+}\hat{\sigma}_{2}^{-}+\hat{\sigma}_{1}^{-}\hat{\sigma}_{2}^{+})$ (with $\hat{\sigma}^{\pm}$ the pseudo-spin operators), which could be tuned. Spin-motion coupling has also be suggested to implement GKP codes [50] for quantum error correction.
@@ Acknowledgements.
Appendix A Condition to neglect the dipole-dipole forces
Assuming that the atoms are in free flight amounts to neglecting the forces induced by the dipole-dipole interaction. This assumption is only valid at short times, when the dipole-dipole force has not yet affected the atomic positions. The time at which this assumption breaks down can be estimated from classical arguments. In the case of Fig. 2, the dominant potential is $\hbar J_{SP}(r)$, which gives rise to a force
$F(r)$ $=-\hbar\frac{dJ_{SP}}{dr}(r)$
$=\frac{3\hbar J_{0}r_{0}^{3}}{r^{4}}$ (4)
where $J_{0}=J_{SP}(r_{0})$. The classical equation of motion for the interatomic distance is
$\mu\frac{d^{2}r}{dt^{2}}=F(r)$ (5)
with $\mu=m/2$ the reduced mass of the two-atom system. If we approximate the force by its value at the initial position $r_{0}$ ($F(r)\simeq F(r_{0})$), we obtain the displacement
$r(t)-r_{0}\simeq\frac{3\hbar J_{0}}{2\mu r_{0}}t^{2}.$ (6)
The time at which this displacement gets on the same order of magnitude as the initial positional disorder ($\Delta r\simeq\sqrt{2}\sigma_{\perp}$) is given by
$t\simeq\sqrt{\frac{2\sqrt{2}\mu r_{0}\sigma_{\perp}}{3\hbar J_{0}}}$ (7)
With the parameters of Fig. 2(b), we find $t\simeq 14$ µs, which is indeed longer than our experimental durations.
Figure 6: Roles of positional disorder and finite Rydberg lifetimes in the damping of a spin-exchange oscillation for the experimental parameters of Fig. 2(c), that is to say $r_{0}=17$ µm, $\theta=90$°, |=|$\ket{\uparrow}=\ket{60S_{1/2},m_{J}=1/2}$ and |=|$\ket{\downarrow}=\ket{60P_{3/2},m_{J}=1/2}$, leading to an average interaction energy $J_{0}=2\pi\times 648$ kHz, $\alpha=3$ and negligible van der Waals interactions. We consider the ideal case of two atoms in the motional ground states of harmonic traps ($\bar{n}_{\perp}=0$), without state preparation errors. (a) Simulated spin-exchange oscillations for atoms in free flight, where the trapping potential is switched off during the exchange, for an initial trapping frequency $\omega_{\perp}=2\pi\times 38$ kHz. Top panel: infinite lifetime ($\Gamma=0$), with the analytical envelope of Eq. (18) as a dotted line). Bottom panel: finite lifetime ($\Gamma\neq 0$). (b) Same in the case of trapped atoms. The analytical envelope is that of Eq. (21). (c) Quality factor as a function of the initial trapping frequency $\omega_{\perp}$. Blue lines indicate the case of atoms in free flight, and orange lines display the case of trapped atoms. Solid lines are simulations with finite lifetimes assuming classical motion; dotted lines are the same simulations with quantum motion; and dashed lines are the case of infinite Rydberg lifetimes. The gray region indicates the inaccessible range of quality factors due to finite Rydberg lifetimes. The points show the parameters used in panels (a,b).
Appendix B A simple analytic model for the damping of a spin exchange oscillation